### v1

# Temporal quantum eraser: Fusion gates with distinguishable photons

Abstract
Linear-optics gates, the enabling tool of photonic quantum information processing, depend on indistinguishable photons, as they harness quantum interference to achieve nonlinear operations. Yet the required indistinguishability is related to the symmetry of the multiphoton wavefunction, and does not necessarily imply identical photons. Here we show for the case of two-photon gates that the ideal gate operation can be retrieved if one can guarantee the exchange symmetry of the input photonic state. Specifically, we employ a temporal quantum eraser to allow fusion gates with sources of modally-impure photons: parametric photon pair generation and single-photon extraction by a single quantum emitter. The ability to lift the requirement for identical photons bears considerable potential in linear-optics quantum information processing.
Introduction.—Photonic quantum information with linear optics utilizes quantum interference in photon detection events to achieve nonlinear operations. Specific detection events at designated output ports of the linear system herald the success of the desired operation. Notable examples include Bell state measurement used for teleportation and for entanglement swapping [1, 2, 3], construction of GHZ states [4, 5], and controlled-NOT gate for linear-optics quantum computing [6, 7]. Of particular interest are fusion gates, first introduced as an efficient method for generating photonic cluster states [8] utilized as a resource for measurement-based quantum computation [9]. More recently, these gates are at the heart of fusion-based quantum computation [10], where they function as destructive entangling measurements, and also serve as a tool in constructing the required small entangled resource states [11].
Figure 1: Useful multiphoton states for linear-optics quantum information processing. The cases in which all the input photons are identical (smaller ellipse), i.e occupying identical modes that propagate in different rails, is only a subset of the useful input states. The broader set comprises of multiphoton states that may involve distinct modes that vary in some degree of freedom, yet exhibit a defined symmetry with respect to exchange of the input rails. In the diagram we chose as an example the polarization discrete degree of freedom, yet our discussion is more general, encompassing multiphoton states with continuous degrees of freedom as well, such as spectro-temporal. The yellow area represents states that are perfectly symmetric to exchange between the input modes (i.e. symmetric to switching the order of the letters in the ket), such as identical photons and symmetric Bell states. The violet area corresponds to states with other exchange symmetries, e.g. the antisymmetric singlet state.
Quantum interference arises from the contribution of indistinguishable paths to the final state, making indistinguishability imperative to the performance of linear-optics quantum information processing [12]. This requirement has promoted the use of identical single photons, which, apart from propagating in distinct rails, share identical properties across all degrees of freedom - polarization, tempo-spectral and spatial modes. Generating such identical photons from different sources presents one of the primary challenges in the field, often involving significant efforts in terms of complexity and resources [12, 13, 14, 15].
However, indistinguishable paths do not necessarily require identical photons; it is sufficient that there is no distinguishing information indicating which input mode resulted in which detection event. In fact, quantum interference is completely dictated by the symmetry of the multiphoton wavefunction with respect to exchange between the inputs. In other words, as illustrated in Figure 1, linear-optics quantum information processing can utilize multiphoton wavefunctions with a defined exchange symmetry. Identical photons, defined by a separable symmetric wavefunction, represent a special case within a broader class of indistinguishable multiphoton states. It should be emphasized that indistinguishable wavefunctions, apart from identical photons, must exhibit some degree of entanglement.
Consider the Hong-Ou-Mandel (HOM) effect, a fundamental two-photon interference phenomenon [16], which has long served as a standard for characterizing photon indistinguishability. This effect involves interfering two photons on a beam splitter (BS), as depicted in Figure 2(a). When the two photons are identical, the paths leading to a single photon at each of the output ports of the BS interfere destructively. This results in photon bunching, where the two photons exit the BS together through either output port. Yet, it can be easily confirmed using the BS input-output relations that the HOM effect is truly a measure of the spatial symmetry of the two-photon state under exchange; a symmetric (antisymmteric) state leads to perfect bunching (antibunching) at the output ports. Therefore, any spatially symmetric or antisymmetric two-photon state, even of distinct internal degrees of freedom, would yield perfect two-photon interference and can hence be utilized reliably in two-photon linear-optics gates.
Figure 2: Hong-Ou-Mandel (HOM) experiment. (a) Two photons are directed towards the input ports of a 50:50 beam splitter. A relative delay between the incoming modes, denoted by $\tau$, is introduced to adjust the overlap between their wavepackets. The HOM effect is observed by measuring the probability of two detection events at different output ports, referred to as the coincidence rate, $R_{c}$. (b) For a symmetric (antisymmetric) two-photon wavefunction, the paths leading to photon antibunching (bunching) at the output ports destructively interfere, resulting in a coincidence rate of zero (one) at zero relative delay. When $\tau$ is set such that the overlap between the modes vanishes, the paths become completely distinguishable and the photons have an equal probability of leaving in either output port, yielding a coincidence rate of half. This leads to the characteristic HOM dip (peak) for a symmetric (antisymmetric) two-photon state, as indicated by the solid (dotted) line.
Note that due to the bosonic nature of photons, antisymmetry in the spatial degree of freedom must be paired with antisymmetry in another degree of freedom. These aspects of two-photon interference have been discussed [17, 18] and demonstrated by using Bell pairs in symmetric and antisymmetric polarization states [19, 20] and hyperentangled photons [21]. A common feature of these demonstrations is that the photons are produced probabilistically in an unheralded manner, thereby limiting their applicability in protocols necessitating on-demand single-photon sources (SPS).
Photon distinguishability, i.e. asymmetry of the photonic state with respect to exchange, can stem from a physical distinction between the modes of the different sources, e.g. different phase matching conditions in non-linear media, or different resonant frequencies of quantum emitters. Yet, even with identical sources, distinguishability may arise from the generation process itself, in which information regarding the mode of the output photon leaks out. When this information is not actively taken into account, the source suffers from mode impurity, resulting in a mixed state for the generated photons, and accordingly, they are at least partially distinguishable.
In this work we focus on the case in which identical copies of such modally-impure sources are used. This means that although the sources are identical, they each produce non-identical, distinguishable photons. In particular, we present two examples in which temporal information associated with the output mode of the photon is carried by additional photons emitted from the same source: parametric photon pair generation in non-linear (NL) media [22, 23, 24, 25], and single-photon extraction by a single quantum emitter [26, 27]. We theoretically demonstrate that by applying a temporal quantum eraser (TQE) on the information from a pair of sources, we can project the two generated photons into a purely symmetric or antisymmetric state. This results in perfect retrieval of the two-photon interference in otherwise distinguishable photons, and enables the performance of two-photon linear optics gates, e.g. fusion gates.
Parametric photon pair generation.—Photon pairs can be stochastically generated in NL optical media through spontaneous parametric down-conversion and spontaneous four-wave mixing, involving materials with $\chi^{(2)}$ and $\chi^{(3)}$ electric susceptibilities, respectively. In these processes, incident pump photons are converted into photon pairs, where each pair consists of a signal photon and an idler photon, generated to separate modes. Ideally, detection of an idler photon in one mode, ensures the presence of the signal photon in the other, thus providing a reliable heralded SPS. Yet, in practical systems, when multiple pairs are produced within the same pump pulse, photon loss and detector inefficiency introduce number impurity to the SPS. To minimize this effect, the generation process is adjusted such that the mean number of photon pairs generated per pulse is significantly lower than one. In turn, the low generation probability can be mitigated by multiplexing several sources, potentially enabling on-demand SPS [28, 29, 30].
Figure 3: Non-linear medium source. Schematic description of a temporal quantum eraser between two $\chi^{(2)}$ sources. A pump pulse is sent to two non-linear media, stochastically generating pairs of signal and idler photons. The idler modes, $\hat{a}_{i}$ and $\hat{b}_{i}$, are mixed on a 50:50 beam splitter and upon two detection events in the output ports within the same pulse, the presence of two signal photons is heralded. The detection pattern is then used to herald the symmetry of the two-photon wavefunction; coalescent (coincident) detection events in $\hat{c}$ and $\hat{d}$ herald a symmetric (antisymmetric) two-photon state at the output.
In order to keep our analysis simple, we examine photons pairs emitted into two guided modes (e.g. two different single-mode waveguides, or opposite directions of the same waveguide), acknowledging that the outcomes are applicable to any choice of separate modes. In what follows, we neglect terms involving multiple pairs. We will address this issue later, when we consider the generation from a pair of sources. Under these assumptions, the state of the signal and idler photons is given by,
[…]
where $\Phi(t_{1},t_{2})$ is determined mostly by the phase-matching condition in the non-linear medium. Assuming the bandwidth of the pump is significantly narrower than the bandwidth of the signal and idler (as is typically the case), energy conservation in the process leads to time-energy entanglement between the generated photons. Hence, temporal information on the signal can be inferred from the detection time of the idler, thus introducing a degree of distinguishability, regardless of the fact that the sources are identical.
Nevertheless, the temporal information carried by idler photons from two identical sources can be ’erased’ by mixing them on a BS, effectively implementing a HOM experiment between the idler photons, as depicted in Figure 3. Consider the state of two photon pairs generated from NL sources with the same pump laser and phase-matching condition,
|$\ket{\psi^{(2)}_{\text{NL}}}=$ $\int dt_{1}dt_{2}dt_{3}dt_{4}\Phi(t_{1},t_{2})\Phi(t_{3},t_{4})\times$ (2)
$a_{i}^{\dagger}(t_{1})a_{s}^{\dagger}(t_{2})b_{i}^{\dagger}(t_{3})b_{s}^{\dagger}(t_{4})\ket{0}$
where $a_{s}$ and $a_{i}$ ($b_{s}$ and $b_{i}$) represent the signal and idler of the first (second) source. Utilizing the the fact that the sources are identical, we can rewrite the state as,
|$\ket{\psi^{(2)}_{\text{NL}}}=$ $\frac{1}{2}\int dt_{1}dt_{2}dt_{3}dt_{4}\Phi(t_{1},t_{2})\Phi(t_{3},t_{4})\times$ (3)
$\sum_{m=0}^{1}\Biggl{[}\Bigl{(}a_{i}^{\dagger}(t_{1})b_{i}^{\dagger}(t_{3})+(-1)^{m}a_{i}^{\dagger}(t_{3})b_{i}^{\dagger}(t_{1})\Bigr{)}\times$
$\Bigl{(}a_{s}^{\dagger}(t_{2})b_{s}^{\dagger}(t_{4})+(-1)^{m}a_{s}^{\dagger}(t_{4})b_{s}^{\dagger}(t_{2})\Bigr{)}\Biggr{]}\ket{0}$
Applying the BS transformation to the idler modes results in,
|$\ket{\psi^{(2)}_{\text{NL}}}=$ $\frac{1}{2}\int dt_{1}dt_{2}dt_{3}dt_{4}\Phi(t_{1},t_{2})\Phi(t_{3},t_{4})\times$ (4)
$\Biggl{[}\Bigl{(}c^{\dagger}(t_{1})c^{\dagger}(t_{3})-d^{\dagger}(t_{1})d^{\dagger}(t_{3})\Bigr{)}\times$
$\Bigl{(}a_{s}^{\dagger}(t_{2})b_{s}^{\dagger}(t_{4})+a_{s}^{\dagger}(t_{4})b_{s}^{\dagger}(t_{2})\Bigr{)}+$
$\Bigl{(}c^{\dagger}(t_{1})d^{\dagger}(t_{3})-c^{\dagger}(t_{3})d^{\dagger}(t_{1})\Bigr{)}\times$
$\Bigl{(}a_{s}^{\dagger}(t_{2})b_{s}^{\dagger}(t_{4})-a_{s}^{\dagger}(t_{4})b_{s}^{\dagger}(t_{2})\Bigr{)}\Biggr{]}\ket{0}$
As indicated by Eq. 3, the symmetry of the idlers is perfectly correlated with that of the signals, where the $m\!=\!0$ and $m\!=\!1$ terms represent the symmetric and antisymmetric states, respectively. Consequently, since the symmetry of a two-photon state dictates the outcome of a HOM experiment, interfering the idlers on a BS enables us to herald the symmetry of the signal photons; coalescent (coincident) detection events in modes $\hat{c}$ and $\hat{d}$ project the signal photons to a symmetric (antisymmetric) state, as described in Eq. 4.
Yet, two detection events in $\hat{c}$ and $\hat{d}$ can also result from either source producing two pairs of signal and idler photons. The relative phase of the pump between the two sources can be adjusted such that the state reads,
|$\ket{\varphi^{(2)}_{\text{NL}}}=\frac{1}{2}$ $\int dt_{1}dt_{2}dt_{3}dt_{4}\Phi(t_{1},t_{2})\Phi(t_{3},t_{4})\times$ (5)
$\Bigl{(}a_{i}^{\dagger}(t_{1})a_{s}^{\dagger}(t_{2})a_{i}^{\dagger}(t_{3})a_{s}^{\dagger}(t_{4})$
$+b_{i}^{\dagger}(t_{1})b_{s}^{\dagger}(t_{2})b_{i}^{\dagger}(t_{3})b_{s}^{\dagger}(t_{4})\Bigr{)}\ket{0}$
Following the BS operation,
|$\ket{\varphi^{(2)}_{\text{NL}}}=$ $\frac{1}{4}\int dt_{1}dt_{2}dt_{3}dt_{4}\Phi(t_{1},t_{2})\Phi(t_{3},t_{4})\times$ (6)
$\Biggl{[}\Bigl{(}c^{\dagger}(t_{1})c^{\dagger}(t_{3})+d^{\dagger}(t_{1})d^{\dagger}(t_{3})\Bigr{)}\times$
$\Bigl{(}a_{s}^{\dagger}(t_{2})a_{s}^{\dagger}(t_{4})+b_{s}^{\dagger}(t_{2})b_{s}^{\dagger}(t_{4})\Bigr{)}+$
$\Bigl{(}c^{\dagger}(t_{1})d^{\dagger}(t_{3})+c^{\dagger}(t_{3})d^{\dagger}(t_{1})\Bigr{)}\times$
$\Bigl{(}a_{s}^{\dagger}(t_{2})a_{s}^{\dagger}(t_{4})-b_{s}^{\dagger}(t_{2})b_{s}^{\dagger}(t_{4})\Bigr{)}\Biggr{]}\ket{0}$
As specified by Eq. 6, coalescent (coincident) detection events at the output ports of the BS herald a symmetric (antisymmetric) two-photon state, in a similar manner to the result of Eq. 4. However, unlike the state in Eq. 4, in this case, both photons occupy either mode $\hat{a}_{s}$ or $\hat{b}_{s}$. Interestingly, since the symmetry of the two-photon wavefunction dictates the outcome of a HOM experiment, perfect two-photon interference of the signal photons persists even when a double pair is generated by either one of the two NL sources. Specifically, the symmetric (antisymetric) state of the signal photons in 6 is characterized by a coincidence rate of zero (one).
Overall, the application of a TQE to idler photons from two NL sources enables us to herald the symmetry of the signal photons, thus retrieving perfect two-photon interference from otherwise partially distinguishable photons. However, the probability of generating double photon pairs from a single source, resulting in two photons in a single waveguide, limits the applicability of this scheme in protocols that encode qubits in single photons. Additionally, in order to achieve an on-demand SPS, heralding on pairs of idler photons would necessitate multiplexing of more sources, as it reduces the average number of generated signal photons compared to heralding on a single idler per source. Nonetheless, the ability to tailor the symmetry of a two-photon state by introducing entanglement may be used as a resource by itself. In the following segment, we introduce a deterministic SPS that avoids these limitations when combined with a TQE.
Single-photon extraction.—Single photon Raman interaction (SPRINT) [31, 27, 32] occurs in a $\Lambda$-type three-level quantum emitter, where the two transitions are exclusively coupled to two orthogonal photonic modes, e.g. two directions of a waveguide (Figure 4(a)). In this process, an incident single photon interacts with a $\Lambda$-system initialized in the bright state associated with the incoming mode. Destructive interference in that mode ’forces’ the quantum emitter to toggle to the corresponding dark state and reflect a photon in the orthogonal mode. As proposed in [26] and demonstrated in [27], when sending a multiple-photon pulse to a quantum emitter in the configuration that leads to SPRINT, the first photon is reflected and the quantum emitter is pushed to its corresponding dark state, thereby transmitting the subsequent photons of the pulse. In this way, a single photon can be extracted from a classical coherent laser pulse with near-unity success probability. Ideally, the only failure mechanism arises from the absence of a photon in the incident coherent pulse, which decreases exponentially with the average number of photons in that pulse, denoted by $\bar{n}$ (Figure 4(b)). Upon an incoming coherent state pulse $f(t)$ in mode $a_{l}$ and a quantum emitter prepared in $g_{r}$, the final state of the photonic field associated with the emitter ending in the respective dark state, $g_{l}$, is given by [26],
|$\ket{\psi^{(1)}_{\Lambda}}=\;$ $-e^{-\bar{n}/2}\sqrt{\bar{n}}\int_{-\infty}^{\infty}dtf(t)a_{r}^{\dagger}(t)\times$ (7)
[…]
As can be seen from the limits of integration in Eq. 7, the transmitted photons carry information about the temporal mode of the extracted single photon. Since the reflected single-photon pulse is guaranteed to end before the quantum emitter becomes transparent to subsequent photons, a detection event in transmission induces a sharp cut-off to the extracted pulse. When tracing over the transmitted field, this time-entanglement lead to a single photon in a mixed state of temporal modes. Therefore, the mode purity, and hence indistinguishability, of extracted photons decreases with the average number of photons in the pulse [33] (Figure 4(b)).
Figure 4: $\Lambda$-type quantum emitter source. (a) The basic principle of SPRINT and single-photon extraction. Two orthogonal modes $\hat{a_{r}}$ and $\hat{a_{l}}$ couple exclusively to transitions |$\ket{g_{l}}\!\leftrightarrow\!\ket{e}$ and |$\ket{g_{r}}\!\leftrightarrow\!\ket{e}$, respectively. A quantum emitter prepared in |$\ket{g_{r}}$ will deterministically reflect (extract) the first photon arriving from a pulse in mode $\hat{a_{l}}$, and undergo Raman transition to the other ground state, thereby becoming transparent to any subsequent photons [27]. (b) The performance a SPS based on extraction of a single photon from a classical coherent pulse. Success probability of extracting exactly one photon (dotted), and HOM coincidence rate with two such sources (solid) as a function of the average number of photons in the incident pulses, $\bar{n}$. As evident, the mode impurity of the generated photons leads to non-zero HOM coincidence rate. (c) Schematic description of a temporal quantum eraser (TQE) between two $\Lambda$-type quantum emitter sources. An incoming coherent pulse, |$\ket{\alpha}$, impinges on a balanced Mach-Zehnder interferometer, where each arm is coupled to a $\Lambda$-system prepared in its left ground state. Circulators divert the extracted single photons in both arms to the different output ports. The parity of the number of detected photons $N_{d}$ in the dark port of the interferometer (which is not dark between the two extraction events) indicates the exchange symmetry of the two-photon state, yielding perfect HOM interference with either zero or unit coincidence rate, as shown in (b).
Incorporating a TQE between two extraction sources allows the temporal information carried by the transmitted fields to be utilized for heralding the symmetry of the two extracted photons, thereby obtaining perfect two-photon interference. Consider the setup depicted in Figure 4(c), in which modes $\hat{a}_{l}$ and $\hat{b}_{l}$ are part of a balanced Mach-Zehnder interferometer. The relative phase between the two arms of the interferometer is set such that, in the absence of quantum emitters, a coherent state at the input interferes constructively (destructively) on $\hat{c}$ ($\hat{d}$), referred to as the bright (dark) port. The extraction procedure begins with the two emitters in their right ground state and an incident coherent pulse with an average photon number $\bar{n}$. We assume that $\bar{n}$ is sufficiently large such that the vacuum component in the two arms of the interferometer is negligible and both photon extractions occur with a unit success probability. Initially, the emitters in their reflective state do not allow the field to be transmitted. Once a single photon has been extracted by only one of the quantum emitters, it becomes transparent and the transmitted field impinging on the output BS can lead to detection events in both the dark and bright port. When the second photon has been extracted and both emitters are transparent, the transmitted fields interfere such that photons are detected only on the bright port.
By adopting Eq. 7 for both arms of the Mach-Zehnder, applying the appropriate BS transformation and rearranging the terms, the final state of the field, associated with both quantum emitters in their respective dark state, can be written as,
|$\ket{\psi^{(2)}_{\Lambda}}=$ $\frac{e^{-\bar{n}/2}\sqrt{\bar{n}}}{2}\int_{-\infty}^{\infty}dt\int_{t}^{\infty}dt^{\prime}f(t)f(t^{\prime})\times$ (8)
$\text{exp}\biggl{(}i\sqrt{\bar{n}}\int_{t^{\prime}}^{\infty}d\tau f(\tau)c^{\dagger}(\tau)\biggl{)}e^{iZ_{c}}\times$
$\sum_{m=0}^{1}\Bigl{(}a_{r}^{\dagger}(t)b_{r}^{\dagger}(t^{\prime})+(-1)^{m}a_{r}^{\dagger}(t^{\prime})b_{r}^{\dagger}(t)\Bigr{)}\times$
$\Bigl{(}e^{Z_{d}}+(-1)^{m}e^{-Z_{d}}\Bigr{)}\ket{0}$
where we define the creation operators for the bright and dark port, acting between the first extraction event, at time $t$, and the second extraction event, at time $t^{\prime}$,
[…]
As evident from Eq. 8, a detection event in the bright port guarantees that at least one of the photons was extracted at some earlier time but it does not provide information on the time ordering between the two extracted photons. In the dark port, detection events can only occur within the timeframe between the two extractions. Since $\bigl{[}e^{Z_{d}}+(-1)^{m}e^{-Z_{d}}\bigr{]}$ for $m=0$ ($m=1$) comprises only of even (odd) powers of creation operator $Z_{d}$, the parity of detection events in the dark port collapses the outgoing state of the extracted photons into two possibilities; an even (odd) number of detection events heralds a temporal symmetric (antisymmetric) two-photon state. Alternatively, the extracted photons can be utilized to herald the generation of a cat state; bunching (antibunching) observed in a HOM experiment between the extracted photons indicates the preparation of an even (odd) cat state in the dark port.
To summarize, employing a TQE in the transmitted fields of two $\Lambda$-type sources allows us to herald the symmetry of the extracted photons, thereby recovering perfect two-photon interference from partially distinguishable photons. Unlike NL sources, since extraction is practically deterministic for $\bar{n}\!\gg\!1$, the introduction of the TQE does not reduce the success probability for photon generation. Moreover, in an ideal setup, only a single photon is extracted from each source, avoiding number impurity concerns. However, it is important to note that the presence of photon loss in realistic systems can hinder an accurate classification of the symmetry of the two-photon state, as the parity of detection events at the dark port is sensitive to such losses.
Fusion gates.—The TQE enables the retrieval of ideal two-photon interference from a pair of imperfect SPS by decomposing the two-photon wavefunction into its symmetric and antisymmetric components. In turn, this allows us to reliably perform fusion gates with such sources, where our knowledge of the symmetry of the two-photon state is utilized to correctly interpret the destructive measurement outcome.
Figure 5: Fusion gates. Schematic description of type-I and type-II fusion gates between two path-encoded photonic qubits, defined in modes $\hat{a}_{0,1}$ and $\hat{b}_{0,1}$, each maximally entangled with a distinct set of qubits, $A$ and $B$, respectively. A type-I fusion gate is implemented by mixing modes $\hat{a}_{1}$ and $\hat{b}_{1}$ on a BS and registering photon detections at its output ports $\hat{d}_{1}^{1,2}$. The success of the gate is heralded by a single detection event in one of these ports, which results in a maximally entangled state of a new qubit, defined in $\hat{a}_{0}$ and $\hat{b}_{0}$, and qubit sets $A$ and $B$. In a type-II fusion, modes $\hat{a}_{0}$ and $\hat{b}_{0}$ are also mixed on a BS and the detection pattern in all output ports is recorded. A successful type-II fusion is heralded by two detection events in the output ports of different BS, which lead to a maximally entangled state of qubit sets $A$ and $B$. The symmetry of the two-photon state in modes $\hat{a}_{0,1}$ and $\hat{b}_{0,1}$, heralded by a TQE in the generation process of the photons, is used to correctly interpret the outcomes of these fusion gates (see Table 1).
As depicted in Figure 5, consider two path-encoded photonic qubits, defined in modes $\hat{a}_{0,1}$ and $\hat{b}_{0,1}$, where each qubit is maximally entangled with a distinct set of qubits, $A$ and $B$, respectively. Applying a TQE in the generation process of photons $\hat{a}$ and $\hat{b}$ heralds the temporal symmetry of the two-photon state. Hence, the initial state can be written as,
|$\ket{\psi_{m}}=$ $\frac{1}{4}\int dt_{1}dt_{2}\varphi_{m}(t_{1},t_{2})\times$ (10)
(|$\Bigl{(}\ket{A_{0}}a_{0}^{\dagger}(t_{1})+\ket{A_{1}}a_{1}^{\dagger}(t_{1})\Bigr{)}\times$
(|$\Bigl{(}\ket{B_{0}}b_{0}^{\dagger}(t_{2})+\ket{B_{1}}b_{1}^{\dagger}(t_{2})\Bigr{)}\ket{0}$
where $m=0$ ($m=1$) represents the symmetric (antisymmetric) wavefunction of photons $\hat{a}$ and $\hat{b}$. A type-I fusion gate involves mixing modes $\hat{a}_{1}$ and $\hat{b}_{1}$ on a BS and placing a detector at each output port. In a type-II fusion gate, modes $\hat{a}_{0}$ and $\hat{b}_{0}$ are also mixed on a BS and the detection pattern of all four detectors is recorded. We denote the detection modes by $\hat{d}^{j}_{k}$, where $k\!\in\!\{0,1\}$ represents the BS mixing modes $\hat{a}_{k}$ and $\hat{b}_{k}$, and $j\!\in\!\{1,2\}$ specifies the two output ports of that BS.
A successful type-I fusion gate is heralded by a single detection event in either $\hat{d}^{1}_{1}$ or $\hat{d}^{2}_{1}$. As evident in Eq. 11, qubit sets A and B, along with a new qubit defined in modes $\hat{a}_{0}$ and $\hat{b}_{0}$, assume a maximally entangled state with a phase that depends on the symmetry of the two-photon state and the location of the detection event.
$d^{j}_{1}(t)\ket{\psi_{m}}=$ $\frac{1}{2\sqrt{2}}\int dt^{\prime}\varphi_{m}(t,t^{\prime})\times$ (11)
(|$\Bigl{(}\ket{A_{1}}\ket{B_{0}}b^{\dagger}_{0}(t^{\prime})+$
$(-1)^{m+j}\ket{A_{0}}\ket{B_{1}}a^{\dagger}_{0}(t^{\prime})\Bigr{)}\ket{0}$
The success of a type-II fusion gate is signaled by a pair of detection events, each at the output of a different BS. As indicated by Eq. 12, the final state of qubit sets $A$ and $B$ is maximally entangled, where the relative phase in the superposition is determined by the detection pattern and the exchange symmetry of the incident two-photon state.
$d^{i}_{0}(t)d^{j}_{1}(t^{\prime})\ket{\psi_{m}}=$ $\frac{(-1)^{j}}{4}\varphi_{m}(t,t^{\prime})\times$ (12)
(|$\Bigl{(}\ket{A_{0}}\ket{B_{1}}+$
$(-1)^{m+i+j}\ket{A_{1}}\ket{B_{0}}\Bigr{)}$
In contrast, as evident in Eq. 13, two detection events at the outputs of the same BS collapse the state of qubit sets $A$ and $B$ to a product state. This results in a failure of the fusion, either type-I when $k\!=\!1$ or type-II for any value of $k$.
$d^{i}_{k}(t)d^{j}_{k}(t^{\prime})\ket{\psi_{m}}=$ $\frac{(-1)^{j}}{4}\varphi_{m}(t,t^{\prime})\times$ (13)
$\Bigl{(}1+(-1)^{m+i+j}\Bigr{)}\ket{A_{k}}\ket{B_{k}}$
Fusion Detection events Symmetric (m=0) Antisymmetric (m=1)
Type-I one event: $d^{j}_{1}$ |$\ket{A_{0},B_{1},0}\!+\!(-1)^{j}\ket{A_{1},B_{0},1}$ |$\ket{A_{0},B_{1},0}\!-\!(-1)^{j}\ket{A_{1},B_{0},1}$
Type-II two events: $d^{i}_{0}d^{j}_{1}$ |$\ket{A_{0},B_{1}}\!+\!(-1)^{i+j}\ket{A_{1},B_{0}}$ |$\ket{A_{0},B_{1}}\!-\!(-1)^{i+j}\ket{A_{1},B_{0}}$
Failure two events: $d^{i}_{k}d^{j}_{k}$ |$\ket{A_{k}}\ket{B_{k}}$ ($i\!=\!j$) |$\ket{A_{k}}\ket{B_{k}}$ ($i\!\neq\!j$)
Table 1: Fusion gates outcomes presented up to normalization and a global sign. A single detection event recorded in either $\hat{d}^{1}_{1}$ or $\hat{d}^{2}_{1}$ indicates a successful type-I fusion. This gate results in a maximally entangled state between qubit sets $A$ and $B$, along with a new qubit whose states |$\ket{0}$ and |$\ket{1}$ are defined by modes $\hat{a}_{0}$ and $\hat{b}_{0}$, respectively. A successful type-II fusion gate involves two detection events at the output ports of different BS, i.e $d^{i}_{0}$ and $d^{j}_{1}$, resulting in a maximally entangled state for qubit sets $A$ and $B$. The successful outcomes in both fusion gates depend on the location of the detection events as well as on the symmetry of the incident two-photon state, as indicated by the sign of the superposition. Failure of the fusion gates, leading to a separable state for qubit sets $A$ and $B$, occurs when two detection events are recorded at the output ports of the same BS.
Overall, the different outcomes of the fusion gates are summarized in Table 1. While the success probability of these gates remains the well-known $1/2$, employing a TQE and heralding on the symmetry of the two-photon state eliminates the infidelity arising from photon distinguishability. In comparison, implementing a fusion gate with photons that are not completely indistinguishable, i.e. their two-photon wavefunction is not purely symmetric or antisymmetric, results in mixing output states with different phases, given by the value of $m$ in Eq. 11 and 12, which leads to a lower degree of entanglement and, hence, gate infidelity.
Discussion.—We have shown that two-photon linear-optics gates rely on the exchange symmetry of the input two-photon wavefunction, and that identical photons (which indeed fall into the category of a symmetric wavefunction) are only a subset of the useful photonic states. Accordingly, even with sources that inherently produce photons in distinct modes due to temporal information leakage during the generation process, perfect two-photon interference can still be achieved from pairs of such sources. This is accomplished by applying a TQE on this information, leading to the heralded symmetrization or antisymmetrization of the spatial part of the two-photon wavefunction, accompanied by a symmetric or antisymmetric joint temporal profile, respectively. As both TQE outcomes exhibit unit visibility in two-photon interference, this approach facilitates a high fidelity implementation of two-photon linear-optics gates, such as fusion gates, with SPS emitting partially distinguishable photons. Furthermore, in $\Lambda$-type sources, the TQE can be used in reverse to deterministically generate cat states.
The main drawback of implementing the TQE scheme between a pair of NL sources is the introduction of number impurity to the output state as a result of double photon pairs generated from one source. This limitation restricts the utilization of the two-photon state in protocols that require at most a single photon per rail. However, two-photon states with defined symmetry, even with multiple photons in each rail, can still be valuable in linear-optics gates. For example, two-photon NOON states have been shown to boost fusion gates [34, 35].
On the other hand, in $\Lambda$-type sources, the primary technical challenge is the susceptibility to photon loss and imperfect single-photon detection efficiency, which can mislead the parity measurement of the number of photons at the dark port of the interferometer. That being said, drastically reducing loss and increasing photon detection efficiency is a major effort in the entire field of photonic quantum computation, which involves, for example, the development of highly efficient on-chip superconducting nanowire single-photon detectors [36, 37]. Analysis of the practical applicability of the TQE is therefore very system-dependant.
The proposed scheme focuses on symmetrization of the two-photon wavefunction generated by a pair of SPS. Generalizing this method to more than a pair of photons falls beyond the scope of this paper; while a general two-photon wavefunction can be straightforwardly decomposed into its symmetric and antisymmetric part by adding or subtracting the two permutations, the process becomes more complex for $n\!>\!2$, involving additional permutation eigenstates. Nevertheless, it may be possible to construct a more intricate, and potentially probabilistic, TQE tailored to meet the specific requirements linear-optics operations involving more than two photons.
Acknowledgements.—We acknowledge support from the Israeli Science Foundation, the Binational Science Foundation, and the Minerva Foundation.